# Consistent energy-momentum trace couplings of fluids

### v1

@@ Abstract
Gravitational models with non-minimal couplings involving the trace of the energy-momentum tensor have become increasingly popular. The idea of coupling the trace of the matter tensor to the geometry can be applied to various matter models, including relativistic perfect fluids. However, it is well-known that the variational formulation of perfect fluids involves some technicalities. We carefully derive the field equations including the trace coupling of a perfect fluid using two different approaches, namely, that given by Brown using Lagrange multipliers, and that given by Schutz using velocity potentials. We show that previous results involving such trace couplings do not match the results presented here. We demonstrate that our fluid’s equations of motion are consistent with the gravitational field equations. Moreover, we present a simple on-shell argument which further supports the correctness of our results.
@@ 1 Introduction
Recent advances in cosmology have revealed that new physics beyond general relativity is necessary to explain observed phenomena. There is a plethora of modified theories of gravity which include alternative geometries with torsion and non-metricity, higher-order theories, and scalar and matter couplings [2]. Non-minimal matter couplings first appeared in the work of Bertolami et al. [3], where curvature is coupled with Lagrangian density of matter.
In 2011 a new idea was proposed in [1] which suggested to couple the Ricci scalar to the trace of the energy-momentum tensor. This has been known as $f(R,T)$ gravity and became hugely influential with over 2300 citations to date, according to INSPIRE. The field equations are correctly stated in [1] in terms of a function that depends on the variation of matter Lagrangian, see also [4]. However, when applied to a perfect fluid, the situation is rather delicate when using $\rho\sqrt{-g}$ or $p\sqrt{g}$ as the fluid Lagrangian. Specifically, one has to be careful with the derived quantity, either the density or the pressure, in order to arrive at the correct field equations when non-minimal couplings are considered.
When working with relativistic perfect fluids, one notices that there is no unique definition of the matter Lagrangian. Several approaches have been shown to lead to the expected energy-momentum tensor of the fluid. The energy-momentum tensor of a relativistic perfect fluid is characterized by the unit four-velocity vector of the flow $U^{\nu}$, the energy density $\rho$ and pressure $p$, and is given by
@@
where $g_{\mu\nu}$ is the metric tensor, see for example [5, 6].
@@
On-shell, meaning that all equations of motion have been taken into account, one does indeed find that $\mathcal{L}_{\rm fluid}=-\rho\sqrt{-g}=p\sqrt{-g}$. However, in a variational approach, one must start with the complete off-shell Lagrangian which will lead to the correct field equations and the fluid’s equations of motion. Only then, can one use the equations of motion in the Lagrangian to get the on-shell expression. A detailed account on the variational approach to relativistic perfect fluids is given in [7, 8]. The consideration of the off-shell Lagrangian for the perfect fluid becomes particularly important when one derives the fluid’s equations of motion. In $f(R,T)$ gravity, the variation of the trace of the energy-momentum tensor introduces new contributions that, as will be shown, modify the expressions for the temperature, number density, and chemical potential. This allows us to define the temperature contribution of the non-minimal coupling, for example, an important physical quantity not previously discussed.
@@
In [7], Brown uses an action functional that depends on Lagrange multipliers and Lagrange coordinates. It introduces Lagrangian coordinates $\alpha^{A}$ that define the flow lines of the fluid on a space-like hypersurface. The fluid’s energy density is given as a function of the particle number density $n$ and entropy $s$, while the pressure $p$ is a derived quantity from the density and its derivative. The Lagrangian is then defined as a function of the aforementioned parameters along with particle number flux $J^{\nu}$ and additional spacetime scalars. This formalism provides a Lagrangian that explicitly depends on the independent parameters which, although cumbersome, allows us to account for all the variational effects.
@@
In an earlier paper, Schutz presented the relativistic perfect fluid Lagrangian using a velocity potentials approach in an Eulerian picture [8]. In this picture, one can identify the Lagrangian as $\mathcal{L}_{\rm fluid}=p\sqrt{-g}$, which looks like the on-shell expression. However, the pressure $p$ here is a function of the chemical potential $\upmu$ and entropy $s$. Contrary to Brown’s approach, the density here is a derived quantity from the pressure. The four-velocity vector in this approach is now a function of six potentials and their gradients and thus in the variational approach, one has to carefully account for the variation of each independent potential. These are $\upmu$ and $s$ along with four additional potentials $\{\phi,\beta,\alpha,\theta\}$. We will present the action functional variation in both Brown and Schutz’s approaches to show the analogy between the results in each and to verify our approach using different Lagrangians.
@@
We start with Brown’s Lagrange multipliers picture in Section 2, with the same analysis done in Schutz’s velocity potential picture in Section 3. In each section, we begin by deriving the field equations and equations of motion. We show that the equations lead to the emergence of effective quantities which allow us to derive the on-shell Lagrangian and introduce new thermodynamic quantities. Finally, we derive the conservation equations starting first from the field equations and then from the fluid equation and demonstrate their equivalence. We shall then see in Section 4 that the results of both pictures lead to the same implications. Specifically, one finds that the non-minimal coupling in $f(R,T)$ merely affects the equation of state when models linear in $T$ are considered.
2 Brown’s Lagrange multipliers approach
@@
  • $n$ is the particle number density
  • $s$ is entropy density per particle
  • $\rho(n,s)$ is the energy density of the matter fluid, a function of $n$ and $s$
@@
    $J^{\mu}=\sqrt{-g}nU^{\mu}\,,\qquad|J|=\sqrt{-g_{\mu\nu}J^{\mu}J^{\nu}}\,,\qquad n=|J|/\sqrt{-g}\,$~, (2.2)
    and $U^{\mu}$ is the fluid’s $4$-velocity satisfying $U_{\mu}U^{\mu}=-1$.
  • $\varphi$, $\theta$, and $\beta_{A}$ are all Lagrange multipliers with $A$ taking the values $1,2,3$, and $\alpha^{A}$ are the Lagrangian coordinates of the fluid.
@@
Now, we can obtain the field equations and equations of motion by applying the variational principle on the action, using the Lagrangian density given by Eq. (2.1). The independent variables which we vary with respect to are:
@@ 2.1 Gravitational field equations
where $f$ is understood to be $f(R,T)$. From this point onward, we also write $f_{R}=\partial f/\partial R$ and likewise $f_{T}=\partial f/\partial T$. The variation of the square root of the determinate of the metric is $\delta\sqrt{-g}=-\frac{1}{2}\sqrt{-g}g_{\mu\nu}\delta g^{\mu\nu}$. The expression of $\delta R$ is well-known and important in the derivation of the field equations of $f(R)$ gravity. It suffices to state the result directly
@@
Variation of $T$
As this calculation has not yet been performed in Brown’s fluid approach, we provide a detailed derivation here. We start from the well-known equation for the energy-momentum tensor for a relativistic perfect fluid
@@
We now vary the trace of the energy-momentum tensor and recall that the energy density is a function of the particle number density and entropy $\rho=\rho(n,s)$,
@@
We note that the expressions in the square brackets are the partial derivatives of the trace with respect to $n$ and $s$
@@
Notice that while the entropy $s$ is an independent variable, the particle number density $n$ is not. The variation $\delta n$ can be further expanded in terms of the independent variables (see detailed derivation in Appendix A.1)
@@
We see that the variation of the trace of energy-momentum tensor not only yields variations with respect to the metric $\delta g^{\mu\nu}$, but also variations of the particle number flux $\delta J^{\mu}$ and entropy $\delta s$. The appearance of entropy contribution in the action variation is particularly interesting as this does not appear before in [1].
Variation of $\mathcal{L}_{\rm fluid}$
@@
We have already derived $\delta\sqrt{-g}$ and $\delta\rho(n,s)$ in terms of the independent variables. We also notice that the last term is written in terms of independent variables and their derivatives only. By varying each term, integrating by parts, and using the definition of $J^{\mu}$ from Eq. (2.2), we find the expression for the variation of the energy-momentum Lagrangian density
@@
The detailed derivation of the expression is given in Appendices A.1 and A.2. We now have contributions from the Lagrange multipliers variations, which we will later use to derive the equations of motion.
The full action variation
We are now ready to substitute the terms in Eqs. (2.7), (2.15), and (2.17) into the action variation in Eq. (2.6). This leads to the full action variation which contains all the information for $f(R,T)$ modified gravity model, with a relativistic perfect fluid expressed in Brown’s picture
@@
By setting $\delta S=0$ and recognising that the equations should hold for arbitrary variations, we can extract the equations for the variation of each independent variable. First, we collect the $\delta g^{\mu\nu}$ terms to get the field equations
@@
which can be written in a more familiar form
@@
Notice here that unlike general relativity and $f(R)$ gravity, the left-hand side of the equation depends on the fluid matter through the dependence on the trace of the energy-momentum tensor. the right-hand side has the familiar $\kappa T_{\mu\nu}$ term along with an extra term that depends on the derivative of $T$ with respect to the particle number density. One is tempted to define
@@
$p_{\rm eff}=p-\frac{1}{2\kappa}f_{T}n\frac{\partial T}{\partial n}\,$, (2.24)
It now appears that the entire effect of the trace coupling has been absorbed into the effective energy-momentum tensor. While this is formally correct, at this point this is misleading as $p_{\rm eff}$ and $T_{\mu\nu}^{\rm eff}$ depend on $f_{T}$ and thus also on the Ricci scalar unless $f_{TR}=0$. This is an expected result for models with non-minimal coupling as one can no longer separate the geometry from the matter on each side of the equation. We will discuss this result further but let us first present the equations of motion derived from the action variation.
@@ 2.2 Equations of motion
This is an expected result as we expect the conservation of the particle number density to hold (Eq. (2.28)). The second condition in Eq. (2.26) restricts the four-velocity direction to constant Lagrange coordinates flow lines. From Eqs. (2.25) and (2.27) we have restriction on the exchange of $\beta_{A}$ and $s$ across the flow.
Furthermore, we can see that the variations with respect to entropy and the 4-vector field lead to
@@
respectively. Now, we get interesting results as both of these equations have terms that emerge from the non-minimal coupling with the trace of the energy momentum tensor. The effects emerging from these modified results will be discussed further in the Section 2.4.
Equation (2.20) and Eqs. (2.25)–(2.30) are the complete set of equations derived from the action. In the following we will interpret the various equations focussing on the effects of the non-minimal coupling. In particular, we will provide a variety of arguments to substantiate our claim that Eq. (2.20) is the correct field equation of $f(R,T)$ gravity when $T$ is the trace of a perfect fluid.
@@ 2.3 On-shell fluid Lagrangian
In the introduction, we claimed that expressing the Lagrangian density as $\mathcal{L}_{\rm fluid}=p\sqrt{-g}$ is only correct on-shell, where the equations of motion have been taken into consideration. One then asks whether we get back the on-shell expression when the equations of motion are incorporated into the Lagrangian. Let us begin recalling with the fluid Lagrangian, Eq. (2.1), and integrating by parts all the derivative terms. This gives
@@
where we note that the final term is a total derivative. By using the equations of motion (2.25)–(2.28), we find that on-shell, the Lagrangian density reduces to
$\mathcal{L}_{\mathrm{fluid}}$ $=-\rho\,\sqrt{-g}\qquad\text{on-shell}\,.$ (2.32)
This is a well-know result and unaffected by the coupling with the energy-momentum scalar, see [7].
Instead, let us start from the equation of motion Eq. (2.30), and contract it with $U^{\mu}$. This gives
@@
The right-hand side is nothing but the expression of the fluid Lagrangian in Brown’s picture. If we recall that the expression that appears on the left-hand side is the effective pressure $p_{\rm eff}$, defined in (2.24). Consequently, this statement is saying that
@@
On-shell, we can also equate the two on-shell expressions and move the extra term of the effective pressure into the density, in which case one finds
$p\,\sqrt{-g}=-\Bigl(\rho-\frac{f_{T}}{2\kappa}n\frac{\partial T}{\partial n}\Bigr)\,\sqrt{-g}=:-\rho_{\rm eff}\,\sqrt{-g}\qquad\text{on-shell}\,.$ (2.36)
This rewriting seems to suggest that there exists a formulation where the energy density of the fluid can be rewritten to absorb the effects of the trace coupling. This is precisely what will happen following Schutz’s velocity potentials approach as will be seen in Eq. (3.13) in Section 3.1.

### v2

@@ Abstract
Gravitational models with nonminimal couplings involving the trace of the energy-momentum tensor have become increasingly popular. The idea of coupling the trace of the matter tensor to the geometry can be applied to various matter models, including relativistic perfect fluids. However, it is well known that the variational formulation of perfect fluids involves some technicalities. We carefully derive the field equations including the trace coupling of a perfect fluid using two different approaches, namely, that given by Brown using Lagrange multipliers and that given by Schutz using velocity potentials. We show that previous results involving such trace couplings do not match the results presented here. We demonstrate that our fluid’s equations of motion are consistent with the gravitational field equations. Moreover, we present a simple on-shell argument which further supports the correctness of our results.
@@ 1 INTRODUCTION
Recent advances in cosmology have revealed that new physics beyond general relativity is necessary to explain observed phenomena. There is a plethora of modified theories of gravity which include alternative geometries with torsion and nonmetricity, higher-order theories, and scalar and matter couplings [1]. Nonminimal matter-curvature couplings are generalizations of $f(R)$ gravity in which an additional matter term is coupled with the Ricci scalar $R$ in the action. The first of these models appeared in the work of Bertolami et al. [2], where curvature is coupled with Lagrangian density of matter in the form of $f(R,L)$. Since then, these models have been found particularly useful in providing an alternative account of the Universe expansion or dark energy, and dark matter[3]. Matter-curvature nonminimal couplings do not conserve the energy-momentum tensor. The covariant derivative of the energy-momentum tensor often results in a nonvanishing term that acts as an additional force term which violates geodesic motion. Thus, these theories have been shown promising in resolving cosmological issues in the weak field regime associated with dark sector physics [4].
In 2011 a new idea was proposed in [5] which suggested to couple the Ricci scalar to the trace of the energy-momentum tensor. This has been known as $f(R,T)$ gravity and became hugely influential with over 2300 citations to date, according to INSPIRE. The trace of the energy-momentum tensor is itself a function of not only the Lagrangian, but also its variations with respect to the metric. The field equations are correctly stated in [5] in terms of a function that depends on the variation of matter Lagrangian; see also [6]. However, when applied to a perfect fluid, the situation is rather delicate when using $\rho\sqrt{-g}$ or $p\sqrt{g}$ as the fluid Lagrangian. Specifically, one has to be careful with the derived quantity, either the density or the pressure, in order to arrive at the correct field equations when nonminimal couplings are considered.
When working with relativistic perfect fluids, one notices that there is no unique definition of the matter Lagrangian. Several approaches have been shown to lead to the expected energy-momentum tensor of the fluid. The energy-momentum tensor of a relativistic perfect fluid is characterized by the unit four-velocity vector of the flow $U^{\nu}$, the energy density $\rho$, and pressure $p$ and is given by
@@
where $g_{\mu\nu}$ is the metric tensor; see for example [7, 8].
@@
On shell, meaning that all equations of motion have been taken into account, one does indeed find that $\mathcal{L}_{\rm fluid}=-\rho\sqrt{-g}=p\sqrt{-g}$. However, the situation is more complicated when nonminimal coupling is present (see [9, 10]). In a variational approach, one must start with the complete off-shell Lagrangian which will lead to the correct field equations and the fluid’s equations of motion. Only then can one use the equations of motion in the Lagrangian to get the on-shell expression. A detailed account on the variational approach to relativistic perfect fluids is given in [11, 12]. The consideration of the off-shell Lagrangian for the perfect fluid becomes particularly important when one derives the fluid’s equations of motion. In $f(R,T)$ gravity, the variation of the trace of the energy-momentum tensor introduces new contributions that, as will be shown, modify the expressions for the temperature, number density, and chemical potential. This allows us to define the temperature contribution of the nonminimal coupling, for example, an important physical quantity not previously discussed.
@@
In [11], Brown uses an action functional that depends on Lagrange multipliers and Lagrange coordinates. It introduces Lagrangian coordinates $\alpha^{A}$ that define the flow lines of the fluid on a spacelike hypersurface. The fluid’s energy density is given as a function of the particle-number density $n$ and entropy $s$, while the pressure $p$ is a derived quantity from the density and its derivative. The Lagrangian is then defined as a function of the aforementioned parameters along with particle-number flux $J^{\nu}$ and additional spacetime scalars. This formalism provides a Lagrangian that explicitly depends on the independent parameters which, although cumbersome, allows us to account for all the variational effects.
@@
In an earlier paper, Schutz presented the relativistic perfect fluid Lagrangian using a velocity potentials approach in an Eulerian picture [12]. In this picture, one can identify the Lagrangian as $\mathcal{L}_{\rm fluid}=p\sqrt{-g}$, which looks like the on-shell expression. However, the pressure $p$ here is a function of the chemical potential $\upmu$ and entropy $s$. Contrary to Brown’s approach, the density here is a derived quantity from the pressure. The four-velocity vector in this approach is now a function of six potentials and their gradients and thus, in the variational approach, one has to carefully account for the variation of each independent potential. These are $\upmu$ and $s$ along with four additional potentials $\{\phi,\beta,\alpha,\theta\}$. We will present the action functional variation in both Brown’s and Schutz’s approaches to show the analogy between the results in each and to verify our approach using different Lagrangians.
@@
We start with Brown’s Lagrange multipliers picture in Sec. 2, with the same analysis done in Schutz’s velocity potential picture in Sec. 3. In each section, we begin by deriving the field equations and equations of motion. We show that the equations lead to the emergence of effective quantities which allow us to derive the on-shell Lagrangian and introduce new thermodynamic quantities. Finally, we derive the conservation equations starting first from the field equations and then from the fluid equation and demonstrate their equivalence. We shall then see in Sec. 4 that the results of both pictures lead to the same implications. Specifically, one finds that the nonminimal coupling in $f(R,T)$ merely affects the equation of state when models linear in $T$ are considered.
2 BROWN’S LAGRANGE MULTIPLIER APPROACH
@@
  • $n$ is the particle-number density;
  • $s$ is entropy density per particle;
  • $\rho(n,s)$ is the energy density of the matter fluid, a function of $n$ and $s$;
@@
    $J^{\mu}=\sqrt{-g}nU^{\mu}\,,\qquad|J|=\sqrt{-g_{\mu\nu}J^{\mu}J^{\nu}}\,,\qquad n=|J|/\sqrt{-g}\,$, (2.2)
    and $U^{\mu}$ is the fluid’s four-velocity satisfying $U_{\mu}U^{\mu}=-1$;
  • $\varphi$, $\theta$, and $\beta_{A}$ are all Lagrange multipliers with $A$ taking the values 1, 2, and 3, and $\alpha^{A}$ are the Lagrangian coordinates of the fluid; and
@@
Now, we can obtain the field equations and equations of motion by applying the variational principle on the action, using the Lagrangian density given by Eq. (2.1). The independent variables which we vary with respect to are
@@ 2.1 Gravitational field equations
where $f$ is understood to be $f(R,T)$. From this point onward, we also write $f_{R}=\partial f/\partial R$ and likewise $f_{T}=\partial f/\partial T$. The variation of the square root of the determinate of the metric is $\delta\sqrt{-g}=-\frac{1}{2}\sqrt{-g}g_{\mu\nu}\delta g^{\mu\nu}$. The expression of $\delta R$ is well known and important in the derivation of the field equations of $f(R)$ gravity. It suffices to state the result directly:
@@
2.1.1 Variation of $T$
As this calculation has not yet been performed in Brown’s fluid approach, we provide a detailed derivation here. We start from the well-known equation for the energy-momentum tensor for a relativistic perfect fluid:
@@ 2.1.1 Variation of $T$
We now vary the trace of the energy-momentum tensor and recall that the energy density is a function of the particle-number density and entropy $\rho=\rho(n,s)$:
@@
We note that the expressions in the square brackets are the partial derivatives of the trace with respect to $s$ and $n$:
@@
Notice that while the entropy $s$ is an independent variable, the particle-number density $n$ is not. The variation $\delta n$ can be further expanded in terms of the independent variables (see detailed derivation in Appendix A.1):
@@
We see that the variation of the trace of energy-momentum tensor yields not only variations with respect to the metric $\delta g^{\mu\nu}$, but also variations of the particle-number flux $\delta J^{\mu}$ and entropy $\delta s$. The appearance of entropy contribution in the action variation is particularly interesting as this does not appear before in [5].
2.1.2 Variation of $\mathcal{L}_{\rm fluid}$
@@
We have already derived $\delta\sqrt{-g}$ and $\delta\rho(n,s)$ in terms of the independent variables. We also notice that the last term is written in terms of independent variables and their derivatives only. By varying each term, integrating by parts, and using the definition of $J^{\mu}$ from Eq. (2.2), we find the expression for the variation of the energy-momentum Lagrangian density:
@@
The detailed derivation of the expression is given in Appendixes A.1 and A.2. We now have contributions from the Lagrange multipliers variations, which we will later use to derive the equations of motion.
2.1.3 The full action variation
We are now ready to substitute the terms in Eqs. (2.7), (2.15), and (2.17) into the action variation in Eq. (2.6). This leads to the full action variation which contains all the information for $f(R,T)$ modified gravity model, with a relativistic perfect fluid expressed in Brown’s picture:
@@ 2.1.3 The full action variation
By setting $\delta S=0$ and recognizing that the equations should hold for arbitrary variations, we can extract the equations for the variation of each independent variable. First, we collect the $\delta g^{\mu\nu}$ terms to get the field equations
@@
which can be written in a more familiar form:
@@
Notice here that unlike general relativity and $f(R)$ gravity, the left-hand side of the equation depends on the fluid matter through the dependence on the trace of the energy-momentum tensor. The right-hand side has the familiar $\kappa T_{\mu\nu}$ term along with an extra term that depends on the derivative of $T$ with respect to the particle-number density. One is tempted to define
@@
$p_{\rm eff}=p-\frac{1}{2\kappa}f_{T}n\frac{\partial T}{\partial n}\,.$ (2.24)
It now appears that the entire effect of the trace coupling has been absorbed into the effective energy-momentum tensor. While this is formally correct, at this point this is misleading as $p_{\rm eff}$ and $T_{\mu\nu}^{\rm eff}$ depend on $f_{T}$ and thus also on the Ricci scalar unless $f_{TR}=0$. This is an expected result for models with nonminimal coupling as one can no longer separate the geometry from the matter on each side of the equation. We will discuss this result further but let us first present the equations of motion derived from the action variation.
@@ 2.2 Equations of motion
This is an expected result as we expect the conservation of the particle-number density to hold [Eq. (2.28)]. The second condition in Eq. (2.26) restricts the four-velocity direction to constant Lagrange coordinates flow lines. From Eqs. (2.25) and (2.27) we have restriction on the exchange of $\beta_{A}$ and $s$ across the flow.
Furthermore, we can see that the variations with respect to entropy and the four-vector field lead to
@@
respectively. Now, we get interesting results as both of these equations have terms that emerge from the nonminimal coupling with the trace of the energy-momentum tensor. The effects emerging from these modified results will be discussed further in the Sec. 2.4.
Equations (2.20) and (2.25)-(2.30) are the complete set of equations derived from the action. In the following we will interpret the various equations focussing on the effects of the nonminimal coupling. In particular, we will provide a variety of arguments to substantiate our claim that Eq. (2.20) is the correct field equation of $f(R,T)$ gravity when $T$ is the trace of a perfect fluid.
@@ 2.3 On-shell fluid Lagrangian
In the introduction, we claimed that expressing the Lagrangian density as $\mathcal{L}_{\rm fluid}=p\sqrt{-g}$ is only correct on shell, where the equations of motion have been taken into consideration. One then asks whether we get back the on-shell expression when the equations of motion are incorporated into the Lagrangian. Let us begin recalling with the fluid Lagrangian, Eq. (2.1), and integrating by parts all the derivative terms. This gives
@@
where we note that the final term is a total derivative. By using the equations of motion (2.25)-(2.28), we find that on shell, the Lagrangian density reduces to
$\mathcal{L}_{\mathrm{fluid}}$ $=-\rho\sqrt{-g}\qquad\text{on shell}\,.$ (2.32)
This is a well known result and unaffected by the coupling with the energy-momentum scalar; see [11].
Instead, let us start from the equation of motion (2.30) and contract it with $U^{\mu}$. This gives
@@
The right-hand side is nothing but the expression of the fluid Lagrangian in Brown’s picture. If we recall that the expression that appears on the left-hand side is the effective pressure $p_{\rm eff}$, defined in (2.24), consequently, this statement is saying that
@@
On shell, we can also equate the two on-shell expressions and move the extra term of the effective pressure into the density, in which case one finds
$p\,\sqrt{-g}=-\Bigl(\rho-\frac{f_{T}}{2\kappa}n\frac{\partial T}{\partial n}\Bigr)\,\sqrt{-g}=:-\rho_{\rm eff}\sqrt{-g}\qquad\text{on shell}\,.$ (2.36)
This rewriting seems to suggest that there exists a formulation where the energy density of the fluid can be rewritten to absorb the effects of the trace coupling. This is precisely what will happen following Schutz’s velocity potentials approach as will be seen in Eq. (3.13) in Sec. 3.1.